OpenStudy (anonymous):

I am unable to solve the two simultaneous equations: 0.991=0.0266x+0.0835y, 0.325=0.0499x+0.0077y. Could you help?

OpenStudy (anonymous):

Use the fisrt equation to find y=f(x) on replace the y's expression found in the second equation to compute x and you'll find y (I have x=0.523...)

OpenStudy (anonymous):

if i were doing it i would clear the decimals and write \[266x+835y=9110\] \[499x+77y=3250\]

OpenStudy (anonymous):

then i would go slap the math teacher for giving me such annoying numbers to work with. then i would use some computer algebra system to solve

OpenStudy (anonymous):

Ha ha...:-)

OpenStudy (anonymous):

well not that hard when you ask the machine. i got \[x=\frac{ 1950680}{396183} \]

OpenStudy (anonymous):

how much arithmetic do you wan to do for a problem like this?

OpenStudy (anonymous):

The thought police will be after u for suggesting an assault on apoor math teacher...

OpenStudy (anonymous):

ok a mild slap with a kleenex. but really, who would even want to know such a thing. if this problem arose naturally from some actual data, then fine. get out the software and find your answer.

OpenStudy (anonymous):

Either u or olive is a bit out...

OpenStudy (anonymous):

i used a machine. i have no stake in this game

OpenStudy (anonymous):

I guess olive used her machine wrong, I'm not going to check it though.

OpenStudy (anonymous):

She got the principle right, that's the main thing...

OpenStudy (anonymous):

Thanks for your help. By the way, this is chemistry...determining unknown concentrations.

OpenStudy (anonymous):

That's a reasonable excuse, no need to slap the chem teacher, then:-)